A satellite S is moving in an elliptical orbit around the earth The mass of the satellite is very small compared to the mass of the earth.
 
(a) The acceleration of S is always directed towards the centre of the earth.
(b) The angular momentum of S about the centre of the earth changes in direction, but its magnitude remains constant.
(c) The total mechanical energy of S varies periodically with time.
(d) The linear momentum of S remains constant in magnitude.

Grade:6

1 Answers

Deepak Patra
9 years ago

(a,b,) For r > R, the gravitational field is F = GM / r2

? F1 = GM/r21 and F2 GM / r22 ? F1/F2 = r22 / r21

For r < R, the gravitational field F = GM/R3 x r

? F1 = GM/R3 x r1 and F2 = GM/R3 x r2

? F1/F2 = r1 / r2